# Adaptively Bypassing Vision Transformer Blocks for Efficient Visual Tracking

### v1

@@ 4.3 State-of-the-art Comparisons
In terms of tracking performance, our ABTrack-AViT surpasses all other real-time trackers across all metrics, while ABTrack-DeiT secures the second position in most metrics on the four test benchmarks. For instance, on TrackingNet, ABTrack-AViT and ABTrack-DeiT take the first and second place respectively, demonstrating superior tracking performance compared to MixFormerV2 by more than 2% across all three performance metrics. Regarding speed, our trackers demonstrate the fastest speeds across all three platforms, except for ABTrack-AViT on AGX, which is slightly lower than HiT. Notably, only ABTrack-DeiT and ABTrack-ViT achieve GPU speeds exceeding 200 FPS. Although inferior to most non-real-time trackers in tracking performance, our trackers’ speeds far exceed theirs. For example, all our trackers are over 7 times faster than the leading tracker ARTrack Wei et al. (2023a) on all three speeds. In comparison to the tracker with the most similar tracking performance, TrSiam Wang et al. (2021a), our ABTrack-AViT achieves over 3 times the GPU speed, over 7 times the CPU speed, and over 5 times the AGX speed. These outcomes confirm the effectiveness of our methodology and its state-of-the-art performance.
Table 2: Comparison with real-time trackers on VOT-ST2021 Kristan et al. (2021).
          Method         EAO     Accuracy   Robustness
ABTrack-AViT 0.277 0.449 0.742
ABTrack-DeiT 0.279 0.445 0.752
ABTrack-ViT 0.235 0.428 0.658
MixFormerV2  Cui et al. (2023) 0.234 0.443 0.637
HiT Kang et al. (2023) 0.232 0.439 0.643
FEAR Borsuk et al. (2022) 0.250 0.436 0.655
LightTrack Yan et al. (2021b) 0.225 0.391 0.641
E.T.Track Blatter et al. (2021) 0.224 0.372 0.631
@@
The VOT-ST challenge series is widely recognized in the tracking community, making it a standard reference for tracking algorithm evaluation. We additionally conducted a comparison with five real-time trackers on the VOT-ST2021 benchmark Kristan et al. (2021), and the results are presented in Table 2. Our ABTrack-AViT and ABTrack-DeiT achieve the best results compared to the other real-time trackers, achieving Expected Average Overlap (EAO), Accuracy, and Robustness of 0.277, 0.449, and 0.742 for the former and 0.279, 0.445, and 0.752 for the latter, respectively.
@@ 4.4 Ablation Study
BDM VTP P AUC FPS (GPU)
ABTrack-AViT 83.8 66.4 181
$\checkmark$ 84.1 0.3$\uparrow$ 67.4 1.0$\uparrow$ 225 24.3$\uparrow$
$\checkmark$ $\checkmark$ 82.3 1.5$\downarrow$ 65.2 1.2$\downarrow$ 248 37.0$\uparrow$
The Impact of the Bypass Decision Module (BDM) and ViT Pruning (VTP). We add the proposed components BDM and VTP into the baseline, i.e., ABTrack-AViT without incorporating the proposed components, one by one to evaluate their effectiveness. The evaluation results on UAV123 are shown in Table 3. As can be seen, the inclusion of BDM leads to a significant increase in speed. Specifically, ABTrack-AViT experiences an increase of 24.3% in GPU speed. Remarkably, the Precision (P) and AUC also increase by 0.3% and 1.0%, respectively, which may be attributed to the enhanced generalization ability of the simplifed model. These substantial improvements of efficiency and the improving of accuracy underscore the effectiveness of BDM. When VTP is further integrated, ABTrack-AViT sees an additional 12.7% increase in GPU speed, with only marginal drops of 1.2% and 0.2% in P and AUC respectively. This improvement in GPU speeds, coupled with minimal compromises in tracking performance, confirms the efficacy of VTP in optimizing tracking efficiency.
Time Overheads of Proposed Components. We delve into the computational time overheads incurred by the proposed components of our method to study their impact on the efficiency of our method. We calculate the time cost of a ViT Block (VB) and the time costs after BDM and VTP being integrated successfully. Note that these time costs, as shown in Table 4, are computed by averaging 10,000 inferences. As can be seen, the time overhead of BDM is about $12\%\approx(0.46-0.41)/0.41$ of the ViT block, and with VTP integrated the time cost of BDM is well offset and remarkably reduce about $9.7\%\approx(0.41-0.37)/0.41$ time cost of VB. These results indicate that bypassing a VB leads to approximately an 88% reduction in the time cost associated with that block. Conversely, if a VB is not bypassed, the time cost is reduced by about 9.7%. These findings underscore the always enhancement in efficiency achieved by the proposed method.
Figure 3: Illustration of tracking results, the number of remaining ViT blocks, and the IOU of ABTrack-DeiT and ABTrack-DeiT ($\tau=\tau_{0}$) across three samples from UAV123 and LaSOT.
Figure 4: The first, second, third, and fourth column shows the original images, the corresponding attention maps of the ABTrack-DeiT*, ABTrack-DeiT, and ABTrack-DeiT ($\tau=\tau_{0}$), respectively. Note that ABTrack-DeiT* denotes the baseline without incorporating the proposed BDM and VTP.
5 Qualitative Results
@@
Module ViT Block (VB) VB + BDM VB + BDM + VTP
time (ms) 0.41 0.46 0.37
Effect of the Block Sparsity Function. To study the effect of the proposed block sparsity function $\tau$ on learning adaptive bypassing ViT blocks, we train ABTrack-DeiT with $\tau$ being substituted by its initial value $\tau_{0}$ for comparison, which is denoted by ABTrack-DeiT ($\tau=\tau_{0}$). The tracking performance of the two methods on UAV123 and GOT-10K are shown in Table 5. It’s evident that ABTrack-DeiT significantly outperforms ABTrack-DeiT ($\tau=\tau_{0}$) on both datasets. Specifically, ABTrack-DeiT achieves gains of 3.6% and 3.1% in Precision (P) and AUC on UAV123, and 3.2%, 3.0%, and 3.8% in AO, $SR_{0.5}$, and $SR_{0.75}$ on GOT-10K respectively. These results strongly support the effectiveness of the proposed block sparsity function for learning adaptively bypassing ViT blocks.
@@ 5.1 Study on the Generality
We seamlessly integrated the proposed approach into various trackers and backbones and conducted evaluations on UAV123 Mueller et al. (2016) to demonstrate its generality.
Application to backbone of various scales. We use three scales of DeiT as backbone, i.e., DeiT-Tiny, DeiT-Small, and DeiT-Base, to study the effectiveness of the proposed method on different scales of backbone. To make sense of each component’s effect, they are integrated into the baselines separately. Note that all super parameters are set to their default values. The P, AUC, and GPU speed on the UAV123 Mueller et al. (2016) dataset are shown in Table 6. As observed, for all these backbones, whether with BDM or both components integrated, there is a substantial improvement in the speed of the baselines, with an increase of over 20.0%. Meanwhile, the reductions in P and AUC are marginal, not exceeding 2.0%. For example, when upon integrating both components into the DeiT-Small baseline, the speed raises by 53.4%, accompanied by a small decrease of 1.5% in P and 1.6% in AUC. It’s noteworthy that when we integrated BDM into DeiT-Tiny, P and AUC actually increase by 0.2% and 0.1%, respectively, compared to the baseline. Additionally, there is a 25.5% increase in speed. Despite performance and speed fluctuations across the backbones, these results evidence the effectiveness of the proposed method across different scales of ViT backbones.
Table 6: Application of our Bypass Decision Module (BDM) and ViT Pruning (VTP) to backbone of various scales, i.e., DeiT-Base, DeiT-Small, and DeiT-Tiny. The evaluation is performed on UAV123.
@@
Table 7: Application of our Bypass Decision Module (BDM) and ViT Pruning (VTP) to SOTA trackers, i.e., DropMAE Wu et al. (2023), GRM Gao et al. (2023) and OSTrack Ye et al. (2022a). The evaluation is on UAV123.
@@
Application to state-of-the-art trackers. To show that our approach is general and can be easily integrated into any ViT-based tracking frameworks, we apply our method to three state-of-of-art trackers, i.e., DropMAE Wu et al. (2023), GRM Gao et al. (2023), and OSTrack Ye et al. (2022a). The two components are subsequently incorporated into these trackers for comparison. The super parameters are set to their default values. The P, AUC, and GPU speed on the UAV123 dataset are shown in Table 7. As can be seen, for all these baseline models, whether incorporating BDM or both components, there is a significant enhancement in speed, showcasing an increase of more than 15.0%. Specifically, when BDM is integrated, the GPU speeds of DropMAE, GRM, and OSTrack rise by 18.8%, 16.5%, and 19.6% respectively. Furthermore, with both components integrated, their speeds increase by 33.5% for DropMAE, 30.1% for GRM, and 34.8% for OSTrack. Regarding tracking performance, we observe that the P and AUC of DropMAE and OSTrack only experience a slight decrease, while those of GRM show an unexpected improvement. Specifically, when BDM is integrated, the P and AUC of DropMAE and OSTrack decrease by less than 1.0%, while those of GRM increase by more than 0.5%. Surprisingly, with both components integrated, the P of GRM is raised by 1.4%. These results validate the applicability of our method to enhance the efficiency of existing ViT-based trackers without significantly compromising their tracking performance. Moreover, it may even enhance the baseline’s performance by reducing redundant parameters, thereby improving its generalization ability.
@@
Figure 3 displays the tracking results, the number of remaining ViT blocks, and the IOU of ABTrack-DeiT and ABTrack-DeiT ($\tau=\tau_{0}$) across three samples from UAV123 Mueller et al. (2016) and LaSOT Fan et al. (2018). Vertical color blocks are utilized to indicate whether the corresponding ViT blocks are retained; a ViT block is bypassed if it appears white. The plots illustrate the variation of the IOU with respect to the frame for both methods. It’s evident that compared to ABTrack-DeiT ($\tau=\tau_{0}$), ABTrack-DeiT can retain more ViT blocks when facing challenges like changes in aspect ratio, occlusions, motion blur, and similar objects. Conversely, it retains fewer ViT blocks when these challenges are absent. Additionally, ABTrack-DeiT demonstrates notably higher IOU scores under these challenging conditions, as indicated by the ellipses marked on the plots. These examples qualitatively validate the effectiveness of the proposed method, particularly the proposed block sparsity loss, in adaptively bypassing ViT blocks for efficient visual tracking.
Figure 4 presents visualizations of attention maps for three samples from LaSOT, generated by ABTrack-DeiT*, ABTrack-DeiT, and ABTrack-DeiT ($\tau=\tau_{0}$), respectively. Note that ABTrack-DeiT* denotes the baseline without incorporating the proposed BDM and VTP. As can be seen, ABTrack-DeiT* and ABTrack-DeiT demonstrate very simlar results, both visually focusing more on the targets than ABTrack-DeiT ($\tau=\tau_{0}$) does. This indicates that the proposed block sparsity function can effectively guide the model to adaptively bypass ViT blocks, outperforming the effectiveness of employing a constant block sparsity. Consequently, the simplified model is capable of producing feature representations as effective as those of the original model.

### v2

@@ 4.3 State-of-the-art Comparisons
Table 1: Compared with the state-of-the-art methods on the GOT-10k [44], LaSOT [45], TrackingNet [47] and UAV123 [49]  [^2] Some methods did not report results on certain datasets in their original papers. We have supplemented these results in our table by testing the official code.. The top three real-time results are displayed in Red, blue and green fonts, while the best non-real-time results are in bold font.
Method Source GOT-10k LaSOT TrackingNet UAV123 Speed(FPS)
AO SR0.5 SR0.75 AUC PNorm P AUC PNorm P AUC P GPU CPU AGX
ABTrack-DeiT 64.0 74.2 56.1 61.1 70.1 62.8 78.1 83.2 74.3 64.4 81.8 223 64 65
ABTrack-ViT 61.9 72.2 53.0 59.9 68.8 60.8 77.2 82.5 73.1 64.3 81.9 241 69 68
MixFormerV2 [31] NIPS23 61.9 71.7 51.3 60.6 69.9 60.4 75.8 81.1 70.4 63.7 81.3 167 45 49
HiT [30] ICCV23 62.6 71.2 54.4 60.5 68.3 61.5 77.7 81.9 73.1 63.3 80.5 173 47 59
E.T.Track [51] WACV23 56.6 64.6 42.5 59.1 66.8 60.1 72.5 77.8 69.5 62.3 81.5 56 32 22
FEAR [29] ECCV22 61.9 72.2 52.5 53.5 59.7 54.5 70.2 80.8 71.5 58.9 79.2 119 34 38
TCTrack [52] CVPR22 66.2 75.6 61.0 60.5 69.3 62.4 74.8 79.6 73.3 60.5 80.0 140 45 41
LightTrack [28] CVPR21 61.1 71.0 54.3 53.8 60.5 53.7 72.5 77.8 69.5 59.9 77.6 134 39 36
Non-real-time ARTrack  [25] CVPR23 73.5 82.2 70.9 70.4 79.5 76.6 84.2 88.7 83.5 67.9 85.9 26 9 8
GRM [22] CVPR23 73.4 82.9 70.4 69.9 79.3 75.8 84.0 88.7 83.3 67.7 85.4 53 18 16
CTTrack [24] AAAI23 71.3 80.7 70.3 67.8 77.8 74.0 82.5 87.1 80.3 68.8 89.5 40 13 12
OSTrack [10] ECCV22 70.5 80.9 66.7 69.1 78.7 75.2 81.9 87.2 81.0 67.5 84.7 68 18 19
CSWinTT [13] CVPR22 69.4 78.9 65.4 66.2 75.2 70.9 81.9 86.7 79.5 70.5 90.3 10 3 3
ToMP [12] CVPR22 72.2 83.9 65.5 68.5 79.2 73.5 81.5 86.4 78.9 68.5 87.3 24 8 7
TrSiam [8] CVPR21 67.3 78.7 58.6 62.4 70.5 60.6 78.1 82.9 72.7 67.4 85.3 55 7 10
@@
In terms of tracking performance, our ABTrack-AViT surpasses all other real-time trackers across all metrics, while our ABTrack-DeiT secures the second position in most metrics on the four test benchmarks. For instance, on TrackingNet, ABTrack-AViT and ABTrack-DeiT take the first and second place respectively, demonstrating superior tracking performance compared to MixFormerV2 with more than 2% gain across all three performance metrics. Regarding speed, our trackers demonstrate the fastest speeds across all three platforms, except for ABTrack-AViT on AGX, which is slightly slower than HiT. Although our trackers may demonstrate tracking performance inferior to most non-real-time trackers, their exceptional speed sets them apart. Specifically, only ABTrack-DeiT and ABTrack-ViT achieve GPU speeds exceeding 200 FPS, and all our trackers surpass the leading ARTrack [25] by more than 7 times across all speed metrics. Additionally, compared to TrSiam [8], a tracker with the closest tracking performance, our ABTrack-AViT achieves over 3 times the GPU speed, over 7 times the CPU speed, and over 5 times the AGX speed. These results underscore the effectiveness of our methodology and its state-of-the-art performance.
Table 2: Comparison with real-time trackers on VOT-ST2021 [50].
Method E.T.Track [51] LightTrack [28] FEAR [29] HiT [30] MixFormerV2 [31] ABTrack-ViT ABTrack-DeiT ABTrack-AViT
EAO 0.224 0.225 0.250 0.232 0.234 0.235 0.279 0.277
Accuracy 0.372 0.391 0.436 0.439 0.443 0.428 0.445 0.449
Robustness 0.631 0.641 0.655 0.643 0.637 0.658 0.752 0.742
@@
The VOT-ST challenge series is widely recognized in the tracking community, is a standard for tracking algorithm assessment. We additionally conducted a comparison with five real-time trackers on the VOT-ST2021 benchmark [50], and the results are presented in Table 2. As can be seen, our ABTrack-AViT and ABTrack-DeiT demonstrate superior performance compared to other real-time trackers, achieving Expected Average Overlap (EAO), Accuracy, and Robustness scores of 0.277, 0.449, and 0.742 for ABTrack-AViT, and 0.279, 0.445, and 0.752 for ABTrack-DeiT, respectively.
BDM VTP P AUC FPS FLOPs(G) Params.(M)
ABTrack-AViT 83.8 66.4 181 2.39 8.27
$\checkmark$ 84.1 0.3$\uparrow$ 67.4 1.0$\uparrow$ 225 24.3$\uparrow$ 0.97-2.39 3.52-8.27
$\checkmark$ $\checkmark$ 82.3 1.5$\downarrow$ 65.2 1.2$\downarrow$ 248 37.0$\uparrow$ 0.88-1.88 3.31-6.36
@@ 4.4 Ablation Study
Impact of the Bypass Decision Module (BDM) and the ViT Pruning (VTP). We add the proposed components BDM and VTP into the baseline (i.e., ABTrack-AViT without incorporating the proposed components) one by one to evaluate their effectiveness. The evaluation results on UAV123 are shown in Table 3. As can be seen, the inclusion of BDM leads to a significant increase in speed. Specifically, ABTrack-AViT experiences an increase of 24.3% in GPU speed. For FLOPs and Params., it is evident that the model’s computational load and parameters decrease after the implementation of the BDM. Notably, the Precision (P) and AUC also increase by 0.3% and 1.0%, respectively, which may be attributed to the enhanced generalization ability of the simplified model. These substantial improvements of efficiency and the improvement of accuracy underscore the effectiveness of our BDM. When the VTP is further integrated, ABTrack-AViT sees an additional 12.7% increase in GPU speed, with only marginal drops of 1.2% and 0.2% in P and AUC respectively. This improvement in GPU speeds, coupled with minimal compromises in tracking performance, confirms the efficacy of our VTP in optimizing tracking efficiency.
@@
Module Time(ms)
ViT Block(VB) 0.41
VB+BDM 0.46
VB+BDM+VTP 0.37
@@
Time overheads of proposed components. We delve into the computational time overheads incurred by the proposed components to study their impact on the efficiency of our method. We calculate the time cost of a ViT Block (VB) and the time costs after the BDM and the VTP being integrated successfully. Note that these time costs, as shown in Table 5, are computed by averaging 10,000 inferences to obtain stable measures. As can be seen, the time overhead of the BDM is about $12\%\approx(0.46-0.41)/0.41$ of the ViT block, and with the VTP integrated the time cost of the BDM is well offset and remarkably reduce about $9.7\%\approx(0.41-0.37)/0.41$ time cost of the VB. These results suggest that bypassing a VB leads to approximately an 88% reduction in the time cost associated with that block. Conversely, if a VB is not bypassed, the time cost is still reduced by about 9.7% due to the VTP. These findings underscore the always enhancement in efficiency achieved by the proposed method.
Effect of the block sparsity function. To study the effect of the proposed block sparsity function $\tau$ on learning adaptive bypassing ViT blocks, we train ABTrack-DeiT with $\tau$ being substituted by its initial value $\tau_{0}$ for comparison, which is denoted by ABTrack-DeiT($\tau=\tau_{0}$). The tracking performance of the two methods on UAV123 and GOT-10k are shown in Table 5. It’s evident that ABTrack-DeiT significantly outperforms ABTrack-DeiT($\tau=\tau_{0}$) on both datasets. Specifically, ABTrack-DeiT achieves gains of 3.6% and 3.1% in Precision (P) and AUC on UAV123, and 3.2%, 3.0%, and 3.8% in AO, SR0.5, and SR0.75 on GOT-10k respectively. These results strongly support the effectiveness of the proposed block sparsity function for learning adaptively bypassing ViT blocks.
Table 6: Ablation study of the block sparsity constant $\tau_{0}$ on the P, AUC, and GPU speed of ABTrack on UAV123.
$\tau_{0}$ 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
P 81.4 80.7 80.8 80.9 79.3 79.1 78.8 78.5 77.8
AUC 63.8 62.5 62.6 63.1 62.1 62.4 61.9 61.5 61.1
FPS 223 235 243 266 286 307 315 319 325
Impact of the block sparsity constant. To investigate the impact of the block spasity constant $\tau_{0}$ on performance, we trained ABTrack-DeiT with values of $\tau_{0}$ ranging from 0.1 to 0.9 and evaluated it on the UAV123. The results are shown in Table 6. Overall, with an increase in the block sparsity constant $\tau_{0}$, the accuracy decrease is small, while the speed improvement is substantial. For example, when $\tau_{0}$ increases from 0.1 to 0.9, the P and AUC respectively decrease by 3.6% and 2.7%, but the speed increases by a remarkable 45.7%. This justifies that our BDM is capable of significantly improving efficiency without causing substantial deterioration in tracking performance, which is useful for deploying ViT-based trackers to platforms with limited computing resources. Noticeable, within the 0.2-0.4 range, there is an observed improvement in accuracy, suggesting that choosing an appropriate $\tau_{0}$ can lead to a more favorable trade-off between speed and accuracy.
Table 7: Ablation study on when to initiate the BDM. Only layers following the layer $n_{enf}$ are incorporated with BDM.
$n_{enf}$ 0 1 2 3 4 5 6 7 8 9
P 82.0 81.6  82.9 82.3 82.5 82.8 83.0 82.8 83.4 83.2
AUC 64.2  64.0 65.2 64.7 64.9 65.0 65.2 65.1 65.7 65.5
FPS 287 276 266 253 246 235 221 210 193 181
avg. Layers 7.2 7.6 8.0 8.4 8.8 9.2 9.6 10.0 10.4 10.8
Study on when to initiate the BDMs. To investigate the impact of the timing of initiating BDMs on performance, we trained ABTrack-DeiT with $n_{enf}$ ranging from 0 to 9 and evaluated on UAV123. The results are shown in Table 7. It can be observed that as $n_{enf}$ increases, P and AUC generally exhibit an upward trend, while FPS decreases. At $n_{enf}=2$, both AUC and FPS rank third, with P ranking fourth. However, when compared to the top performance of 83.4% P and 65.7% AUC at $n_{enf}=8$, its P and AUC are only 0.5% lower, while the speed is 73 FPS higher. Although the speed decreases by 10 FPS when compared to $n_{enf}=1$, the increase in P and AUC is 1.3% and 1.2%, respectively. This illustrates a nuanced balance between speed and accuracy, showing that $n_{enf}=2$ provides a good compromise, significantly boosting speed without a substantial sacrifice in performance.
Impact of global and local ranking on ViT pruning. In this section, we analyze the impact of global and local ranking methods on the pruning effectiveness of ViTs. The objective is to understand how these two strategies affect the model’s performance and efficiency. The impact of global and local ranking strategies on ViT pruning was assessed using the UAV123 and GOT-10k datasets. The evaluation results, presented in Table 8, reveal that the local ranking strategy significantly outperforms the global ranking strategy. Specifically, the local ranking strategy achieves a 1.6% and 1.7% improvement over the global ranking strategy in Precision (P) and AUC, respectively, on UAV123. On GOT-10k, it leads to improvements of 1.9% in AO, 2.2% in SR0.5, and 1.8% in SR0.75. These results indicate that the local ranking method is more effective for pruning ViT models, leading to better performance across multiple metrics on both datasets. This superiority is likely due to the ability of local ranking to more accurately identify and retain the most critical features within each layer, whereas global ranking may overlook layer-specific importance, resulting in less optimal pruning decisions.
Table 8: Impact of global and local ranking on ViT pruning.
ranking strategy UAV123 GOT-10K
P AUC AO SR0.5 SR0.75
ABTrack-DeiT global ranking 80.2 62.7 62.1 72.0 54.3
local ranking 81.8 64.4 64.0 74.2 56.1
@@ 4.5 Study on the Generality
To demonstrate the generality of our proposed approach, we integrated it into multiple ViT backbones and state-of-the-art trackers. We then evaluated the performance on the UAV123 dataset [49]. This evaluation highlights the versatility and effectiveness of our method across different models. All hyperparameters were set to their default values.
Application to backbone of various scales. We use three scales of DeiT as backbones, i.e., DeiT-Tiny, DeiT-Small, and DeiT-Base, to study the effectiveness of the proposed method on different scales of backbones. To make sense of each component’s effect, they are integrated into the baselines separately. The experimental results are shown in Table 10. As observed, for all these backbones, whether with BDM or both components integrated, there is a substantial improvement in the GPU speed of the baselines, with an increase of over 20.0%. Meanwhile, the reductions in P and AUC are marginal, not exceeding 2.0%. For example, when integrating both components into the baseline with DeiT-Small as the backbone, the speed raises by 53.4%, accompanied by only marginal decreases of 1.5% in P and 1.6% in AUC. It’s noteworthy that integrating BDM into the baseline with DeiT-Tiny as the backbone results in a 0.2% increase in P and a 0.1% increase in AUC. Additionally, there is a 25.5% increase in speed. These results highlight the adaptability and effectiveness of the proposed method across various scales of ViT backbones, demonstrating its potential to enhance both speed and performance in visual tracking tasks.
Table 9: Application of BDM and VTP to backbone of various scales, i.e., DeiT-Base, DeiT-Small, and DeiT-Tiny.
@@
Table 10: Application of BDM and VTP to SOTA trackers, i.e., DropMAE [23], GRM [22] and OSTrack [10].
@@
Application to state-of-the-art trackers. To show that our approach is general and can be easily integrated into any ViT-based tracking frameworks, we apply our method to three state-of-of-art trackers, i.e., DropMAE [23], GRM [22], and OSTrack [10]. The two components are subsequently incorporated into these trackers for comparison. The experimental results are shown in Table 10. As can be seen, all baseline models experience a substantial increase in GPU speed, whether integrating BDM alone or both components together, with improvements exceeding 15.0%. Particularly, integrating BDM alone enhances the GPU speed of OSTrack by 19.6%. Moreover, with both components integrated, its speed further rises to 34.8%. Regarding tracking performance, we observe that the P and AUC of DropMAE and OSTrack experience only a slight decrease, while GRM unexpectedly shows an improvement. Specifically, when BDM is integrated, the P and AUC of DropMAE and OSTrack decrease by less than 1.0%, while those of GRM increase by more than 0.5%. Surprisingly, with both components integrated, the P of GRM is raised by 1.4%. These results validate the applicability of our method to enhance the efficiency of existing ViT-based trackers without significantly compromising their tracking performance. Moreover, it may even enhance the baseline’s performance by reducing redundant parameters, thereby improving its generalization ability.
Figure 4: Illustration of the number of remaining ViT blocks, the IOU, and the predicted bounding boxes of ABTrack-DeiT and ABTrack-DeiT($\tau=\tau_{0}$) across three samples from LaSOT.
5 Qualitative Results
@@
To make sense of the effectiveness of the proposed block sparsity loss in guiding the model to dynamically bypass ViT blocks according to the tracking task, we visualize in Figure 4 the number of remaining ViT blocks alongside the Intersection over Union (IOU) and the predicted bounding boxes of ABTrack-DeiT and ABTrack-DeiT($\tau=\tau_{0}$) across three samples from the LaSOT dataset [45]. The color blocks in the vertical bars indicate retained ViT blocks, while those white blocks represent bypassed ViT blocks. The plots illustrate the variation of the IOU with respect to the frame for both methods. As oberved, when faced with challenging conditions such as out-of-view targets, distractors, changes in aspect ratio, and rotations, ABTrack-DeiT tends to retain more ViT blocks than ABTrack-DeiT($\tau=\tau_{0}$). This adaptive behavior suggests that the block sparsity loss effectively guides the model to preserve essential information for maintaining tracking accuracy under complex conditions. Conversely, in scenarios where tracking conditions are less demanding and these challenges are absent, ABTrack-DeiT retains fewer ViT blocks than ABTrack-DeiT($\tau=\tau_{0}$). This efficiency-driven adaptation demonstrates the proposed block sparsity loss is able to reduce computational overhead by guiding the model to bypass unnecessary computations when simpler scenarios are encountered. Moreover, ABTrack-DeiT demonstrates notably higher IOU scores under these challenging conditions, as indicated by the ellipses marked on the plots. These examples qualitatively emphasize the effectiveness and benefits of the proposed method, particularly the proposed block sparsity loss, in adaptively bypassing ViT blocks for visual tracking.
Figure 5: Qualitative analysis on five video sequences. These sequences were sourced from LaSOT and UAV123.
Figure 5 presents a qualitative tracking analysis comparing our ABTrack-AViT with four state-of-the-art real-time trackers across five diverse video sequences. As illustrated, our tracker is the only one that consistently tracks the target in all challenging scenarios. These scenarios include scale variation (all examples), pose variation (bird-17, dog-19), similar objects (bird-17, cup1, dog-19), complex motion patterns (boat8), and occlusion (truck1). These results further underscore the effectiveness of our proposed approach.